# Beyond the EPICS: comprehensive Python IOC development with QueueIOC

### v3

@@ 4 Device IOCs and “soft” IOCs with QueueIOC
QueueIOC and s6-epics (the latter to be covered below) have been released, respectively, at https://codeberg.org/CasperVector/queue_iocs and https://codeberg.org/CasperVector/s6-epics; fully open-source editions of Mamba and ihep-pkg have been released, respectively, at https://codeberg.org/CasperVector/mamba-ose and https://codeberg.org/CasperVector/ihep-pkg-ose. QueueIOC itself depends on some patches for caproto; GUIs in the open-source edition of Mamba (including the standalone ones, which are in the mamba_lite subdirectory) depend on some patches for pyqtgraph; the supplementary materials include OPIs which depend on some patches for PyDM; all these currently HEPS-specific patches are available from the open-source edition of ihep-pkg.
Our actual introduction to QueueIOC begins with a workalike of StreamDevice, which is in our eyes a simplest yet most useful EPICS module: the QScanIOC class, in combination with classes like TimedRWPair; an example IOC for this, qscan_b2985, is given for the Keysight B2985 electrometer, shown in Figure 5. As can be seen from the figure, QScanIOC provides friendly encapsulation for periodically polling (“scanning”) of devices, while TimedRWPair (just like the asyn module) encapsulates communication via serial ports or TCP/UDP. Unlike StreamDevice which uses “protocol files”, essentially a still quite limiting domain-specific language (DSL), QScanIOC-based IOCs have native access to Python’s capability to process textual and/or binary data. As these IOCs also have full access to Python’s expressiveness, we can easily abstract boilerplate code, eg. with cmd_map in qscan_b2985, which would be clumsy to do with EPICS databases. A more complex example is the qscan_hfda IOC, which will be detailed in Section 6. Based on the extensive support for communication protocols available with Python (whether in standard libraries or third-party libraries), with QueueIOC it is also easy to write IOCs based on more complex interfaces like JSON, HTTP, modbus etc: eg. the qdet_eiger IOC for the Eiger detector uses the vendor’s protocols based on JSON/HTTP and ZeroMQ. qdet_eiger is based on QDetectorIOC, a QueueIOC-based framework for detector integration presented in an accompanying paper (Zhang et al., 2024), which aims to overcome some architectural limitations of the areaDetector framework while still offering decent performance.
@@
QScanIOC is intended for devices with basically stateless interfaces, whose main event loops only differ in the logic that can be separated into on_init(), on_scan() and on_req(). This is why the main loop in QScanIOC does not need to be explicitly written for each device; for most other devices, explicit main loops are necessary. For training purposes in writing main loops, the qdet_nct IOC for the Tsuji (N)CT counter and the qdet_o974 IOC for the Ortec 974 counter are good examples, both based on QDetectorIOC. QScanIOC is not used as we want them to support areaDetector-like acquisition of a fixed (and possibly infinite) number of frames, a feature unfit for stateless implementations. qdet_nct can automatically detect the number of counter channels on the specified device, and adjust its set of PVs accordingly, without relying on code generators (cf. Section 1). QueueIOC’s automatic adaption to dynamic hardware interfaces (cf. also Figure 8) is based on Python’s type() mechanism for runtime creation of classes; based on this mechanism, we can also create advanced abstractions for variable sets of detector features, similar to those in the ADGenICam and ADEiger IOCs but with much more succinct code (Zhang et al., 2024). As can be seen from Figure 5, IOCs based on QueueIOC, whether for devices with dynamic interfaces or not, are easy to run for operators: unlike st.cmd in EPICS, with these Python IOCs device identifiers (addresses, ports etc) and other tuning parameters can be given on the command line, which is well separated from the IOCs’ source code. Additionally, when necessary it is also quite easy to customise IOC behaviours on deeper levels, without needing to copy entire IOC source files: eg. the file helpers/st_b2985.py shows how to customise the behaviours of qscan_b2985 in two aspects.
@@
Till now all IOCs discussed in this section interact with hardware other than the controlling computer; besides “sequencers” which will be covered in Section 5, pure-software IOCs can also be written with QueueIOC. Our example for pure-software QueueIOC-based IOCs is qioc_s6, a workalike for the procServControl IOC, based on s6-epics, a workalike of procServ (Thompson, 2004). s6-epics itself is based on s6 (https://skarnet.org/software/s6/), a well-designed suite of programs to manage service processes; that latter most importantly (to us and concerning EPICS) supports separate logging for each service with reliable log rotation to avoid exhaustion of disk space. After organising IOCs in the ~/iocBoot convention, we can use administration commands like those in (Liu, Dong and Li, 2022), and a few additional useful commands (Figure 6). As can be seen from the figure, aside from the ability to automatically start specified IOCs upon booting, a currently unique feature of s6-epics is the ability to let both caproto-based IOCs and EPICS IOCs exit gracefully. With qioc_s6, the management of IOCs controlled by s6-epics can be done through the CA protocol; PyDM OPIs (Figure 7) have also been developed for it, available from the supplementary materials.
@@
In EPICS, “sequencers” based on the seq module are used to implement modules like optics, sscan etc; meanwhile, as will be detailed in Section 6, writing seq-based “sequencers” can be error-prone. With QueueIOC the implementation can be much cleaner, by leveraging the expressiveness of Python and following a succinct approach. We would also like to note that as has been hinted in Figure 4(b), a less noticed aspect of “sequencers” is the automated manipulation of PVs in other IOCs through the CA protocol; in our humble opinion this, instead of state transitions, is the real essence of EPICS “sequencers”. Consequently, in QueueIOC although there is a class QSequencerIOC for general state machines, all “sequencers” discussed in this section are actually based on its subclass QMotorSeqIOC. The latter creates state machine with only two states, “down” and “up”, in order to implement a workalike of seq’s “all channels connected & received 1st monitor”. As the name implies, it is mainly used to implement “sequencers” IOCs that (just like those in optics and sscan) manipulate motor IOCs, which we think is a most common application scenario for them. Similar to their EPICS counterparts, all these IOCs expose some motor-like interface, which we have intentionally designed to be unified; PyDM OPIs (Figure 7, available from the supplementary materials) are given for them, along with a Bluesky encapsulation (available as QueueMotor in the open-source edition of Mamba, cf. Section 4).
@@
The simplest “sequencer” IOCs are used to systematically prevent collision (bumping) between motors; they are based on the class QFuncbumperIOCBase, with qbumper_sim (Figure 8) as a simple example IOC. A more complex QFuncbumperIOCBase-based IOC has been deployed at the transmission X-ray microscope beamline (BE) of HEPS, while more similar IOCs are expected to be deployed at quite a few of the 15 beamlines of HEPS Phase I. As can be seen from qbumper_sim, the complicated logic in the main loop of these “sequencers” IOCs are abstracted in the library, so that developers only need to specify the essential information; in the case of anti-bumping, the information is the mathematical constraints and the list of motor PVs. Under the hood, QFuncbumperIOCBase maintains the state machine in accordance with the underlying motor IOCs; based on this, it denies motion requests when the specified constraints would be violated, when any motor correlated with the requested motor (including itself) is already moving, and when the machine is in the “down” state. Depending on the requirements for other QueueIOC-based “sequencer” IOCs, the information that needs to be specified by the developer can be less mathematical; among them are IOCs based on the class QMotorHubIOCBase, used to control motors connected to multiplexer PLCs (“motor hubs”) for motion controllers. An example IOC based on the class is qmhub_b5, which is actually used in attitude tuning (Li et al., 2024) of the Raman spectrometer at the hard X-ray high-resolution spectroscopy beamline (B5) of HEPS, with a simulated variant in the open-source edition of Mamba (cf. also Figure 67). A particularly notable feature of QMotorHubIOCBase is a delay between the latest motion of a motor and the switch from it to any other motor connected to the same multiplexer as this motor, along with another delay after this switch before the newly chosen motor can be moved; both delays are tunable, and are enforced to prevent potential power surges that may damage the devices involved.
@@
Systematic support in QueueIOC for monochromators, a perhaps most widely known application scenario for “sequencer” IOCs, is provided in the class QMonochromatorIOCBase. Based on it, workalikes of the optics IOCs for double-crystal monochromators and high-resolution monochromators are respectively given as qmono_dcm and qmono_hr; a usage example for the latter is given in the supplementary materials. For training purposes, a simple “monochromator” IOC is provided as qmono_sim, which involves two axes and simply uses their sum or difference as the “energy” value. Apart from the basic features, QMonochromatorIOCBase also provides support for automated speed tuning to make the beam change smoothly when the energy value is changed; of course, just like its counterpart in the optics IOC, this class only implements a linear approximation to the ideal behavior. In qmono_dcm and qmono_hr, support is also provided for the user-friendly specification of Miller indices and lattice parameters. This feature, as well as the speed-tuning feature above, are also provided in ways we believe to be most friendly for IOC developers: eg. the latter is enabled by default, and does not need any code to be written by the developer; it can also be opted out by simply overriding the _auto_velo member of QMonochromatorIOCBase. As can be seen from the source code of these IOCs, in our eyes the essence of monochromators IOCs are coordinate transformations between the energy value and the motor positions; calibration is the process of setting up parameters for the transformations before the latter are put into actual use. Therefore QMonochromatorIOCBase can also be used for multiple-to-multiple coordinate transformations: eg. those between the reciprocal and direct spaces in crystallography, where the transformation parameters may be imported from specialised programs like xrayutilities (https://github.com/dkriegner/xrayutilities) and diffcalc (https://github.com/dls-controls/diffcalc). By exposing transformation parameters as PVs, like what is done in the QMonochromatorIOCBase-based IOCs discussed above, the importing process can be automated. Using the PV interface, the parameters can also be exported to other programs: eg. trajectory programs used in fly scans involving monochromators or the reciprocal space, which can achieve vastly more accurate motion behaviours than what is possible with the speed-tuning feature of QMonochromatorIOCBase.
6 Case studies: some laser controllers and monochromators
@@
In Section 3, we noted that QueueIOC attempts to replace most EPICS IOCs in a systematic way, while striving to keep the new IOCs as simple as reasonable. We realise that after the brief tour of functionalities in Sections 45, it may still not be obvious what the unique benefits of QueueIOC are; so in this section, we analyse these benefits in detail by having a close look at some example IOCs. Our first example is the qscan_hfda IOC for CNI PSU-H-FDA and PSU-A-D laser controllers. In this IOC, the ancillary class PyHfda and function make_hfda() (Figure 9) are more notable than the IOC class HfdaIOCBase. PSU-H-FDA (older) and PSU-A-D (newer) share a basic serial-based communication protocol, but the older model does not support state readback (cf. the function read_state() in PyHfda); the state readback provides more than 10 information items, all of which are mapped into PVs by the IOC. The state-changing commands require checksum bytes and produce echo replies, so in case a setpoint item is not available from the state readback, the IOC needs to decide whether to update the value of the corresponding PV upon user writing based on the correctness of the replies to the commands. For this kind of readback-less setpoints (enable_val, as well as current_val and power_val for the older model), the IOC also needs to properly initialise the corresponding hardware settings, so that the initial values of their PVs are correctly reflected. Moreover, the newer model is known to encounter read timeouts if a state-reading command is sent too closely after a state-changing command, so an additional delay should be enforced after the latter. All the above are handled succinctly and cleanly in qscan_hfda: state readback in read_state() based on a call to struct.unpack(); complete correctness check in send_cmd() and read_state() based on checksums and echo replies; the automatic differentiation between the old and new models, as well as the preparation of initial states, in make_hfda(). Done in HfdaIOCBase are the automatic construction of PV lists for different models, the different handling of state changing depending on the model, and the delay after state changing. While these are not impossible in EPICS IOCs, they would be very awkward to implement: what the struct.unpack() does above would be highly bloated with StreamDevice; additionally StreamDevice cannot easily support the echo check and the delay after state changing, and we also do not find seq-based “sequencers” very helpful on this issue. Writing the IOC in C/C++ may be also considered, but the codebase would still be significantly larger, especially considering the support for different versions; these make the development and maintenance cost of such an IOC much higher than our IOC.
@@
What distinguishes QueueIOC from the others is an intentional pursuit of the utmost simplicity (Hoare, 1981), or in other words approaching the complexity lower-bounds (cf. Section 1); what QueueIOC attempts to do, in practice, is to provide succinct yet powerful tools that developers can use to build IOCs that are close to their complexity lower-bounds. Admittedly, with caproto, PCASPy etc, it would not be hard to write workalikes of qscan_hfda that are not much more complex than it. This is because its logic revolves around simple state-changing and state-reading commands, and the abstraction for this pattern, QScanIOC, is easy to reimplement; as has been hinted above, PyHfda and make_hfda() are what really make the IOC simple. There are also examples where the tools provided by QueueIOC significantly contribute to the simplicity of the IOCs: in fact, all IOCs for monochromators, motor anti-bumping and motor multiplexing in Section 5 can be seen as examples for this; it can be easily seen when the reader considers how the same requirements could be implemented with EPICS, caproto, PCASPy etc. Here we also discuss monochromators in detail, as another practical demonstration of how simplicity is achieved in QueueIOC-based IOCs – with not only particular tools, but also reusable design patterns. Writing seq-based “sequencers” is often error-prone, because the state transitions in them can easily become comparable to the “goto hell” C/C++ (Figure 10c), result in hard-to-trace bugs (a few of them are given in the supplementary materials). This is why most “sequencers” in QueueIOC only have the “down” and “up” states (cf. Section 5); the domain-specific business logic is instead encapsulated in the state-transition functions. In monochromator IOCs based on QMonochromatorIOCBase, the functions serve_down(), serve_up(), do_common() etc are the basis of our abstraction of their business logic as coordinate transformations. However, in addition to these shared logic, real-world monochromators need invertible three-way conversions between the Bragg angle $\theta$, the wavelength $\lambda$ and the energy $E$; the conversions can be even more complex for certain monochromators, eg. the high-resolution monochromator (Figure 10a). To do this cleanly, we use functions to encode these conversions: eg. mono_phis() in the qmono_hr IOC (Figure 10b), as well as mono_theta() which is used by qmono_dcm and usable by other simple variants of the double-crystal monochromator.
@@
In comparison with the monochromator IOCs in optics, IOCs like qmono_dcm are not only much shorter, but also much better modularised. The shared logic is contained in the code of QMonochromatorIOCBase, organised cleanly unlike the “goto hell”. The $\theta$/$\lambda$/$E$ conversions are fully encapsulated in functions like mono_phis(), which are implemented as straightforward encodings of graphs like Figure 10(a). The rest cannot be easily abstracted as libraries, but they are still organised cleanly, with a structure intentionally kept consistent across different monochromator IOCs; this can be seen from a comparison between the IOC classes, eg. MonoHrIOC in qmono_hr and MonoDcmIOC in qmono_dcm. The result of the structural differences above is dramatic reduction in the cost of development and maintenance. When writing a new monochromator IOC, the developer often only needs to customise the conversion function and the IOC class, without worrying about the possibility of a small change conflicting with some assumptions implicit in the global state. When debugging, issues can usually be easily traced and resolved thanks to the careful decoupling: eg. the conversion function can isolated and used as a calculator for the variables involved, greatly facilitating tests. The same approach is also followed elsewhere in QueueIOC for non-trivial IOCs: aside from PyHfda in qscan_hfda, another example is QDetectorIOC (Zhang et al., 2024), where C/C++ libraries (conventionally called “SDKs” or software development kits) from vendors are encapsulated thinly into self-made mini-“SDKs”; the latter expose succinct, reusable and adaptable interfaces, which can be tested in isolation and even reused in standalone applications. With measures taken like those above, the IOCs provided by QueueIOC have been made satisfactorily close to their complexity lower-bounds. As has been noted in Section 1, improvements in simplicity coincide with improvements in efficiency; with the consistent pursuit of simplicity throughout QueueIOC, we believe it is able to bring about significant efficiency boosts in the EPICS ecosystem.

### v4

@@ 4 Device IOCs and “soft” IOCs with QueueIOC
Our actual introduction to QueueIOC begins with a workalike of StreamDevice, which is in our eyes a simplest yet most useful EPICS module: the QScanIOC class, in combination with classes like TimedRWPair; an example IOC for this, qscan_b2985, is given for the Keysight B2985 electrometer, shown in Figure 5. As can be seen from the figure, QScanIOC provides friendly encapsulation for periodically polling (“scanning”) of devices, while TimedRWPair (just like the asyn module) encapsulates communication via serial ports or TCP/UDP. Unlike StreamDevice which uses “protocol files”, essentially a still quite limiting domain-specific language (DSL), QScanIOC-based IOCs have native access to Python’s capability to process textual and/or binary data. As these IOCs also have full access to Python’s expressiveness, we can easily abstract repeated code, eg. with cmd_map in qscan_b2985, which would be clumsy to do with EPICS databases. A more complex example is the qscan_hfda IOC, which will be detailed in Section 6. Based on the extensive support for communication protocols available with Python (whether in standard libraries or third-party libraries), with QueueIOC it is also easy to write IOCs based on more complex interfaces like JSON, HTTP, modbus etc: eg. the qdet_eiger IOC for the Eiger detector uses the vendor’s protocols based on JSON/HTTP and ZeroMQ. qdet_eiger is based on QDetectorIOC, a QueueIOC-based framework for detector integration presented in an accompanying paper [19], which aims to overcome some architectural limitations of the areaDetector framework while still offering decent performance.
@@
QScanIOC is intended for devices with basically stateless interfaces, whose main event loops only differ in the logic that can be separated into on_init(), on_scan() and on_req(). This is why the main loop in QScanIOC does not need to be explicitly written for each device; for most other devices, explicit main loops are necessary. For training purposes in writing main loops, the qseq_bestec IOC in Section 6 is an expressive example; the qdet_nct IOC for the Tsuji (N)CT counter and the qdet_o974 IOC for the Ortec 974 counter are also good examples, both based on QDetectorIOC. QScanIOC is not used as we want them to support areaDetector-like acquisition of a fixed (and possibly infinite) number of frames, a feature unfit for stateless implementations. qdet_nct can automatically detect the number of counter channels on the specified device, and adjust its set of PVs accordingly, without relying on code generators (cf. Section 1). QueueIOC’s automatic adaption to variable device interfaces (cf. also Figure 8, the qseq_bestec IOC and the self-made SDKs in Section 6) is based on Python’s mechanisms for runtime creation and using of classes; based on these mechanisms, we can also create advanced abstractions for variable sets of detector features, similar to those in IOCs based on ADGenICam but with much more succinct code [19]. As can be seen from Figure 5, IOCs based on QueueIOC, whether for devices with variable interfaces or not, are easy to run for operators: unlike st.cmd in EPICS, with these Python IOCs device identifiers (addresses, ports etc) and other tuning parameters can be given on the command line, which is well separated from the IOCs’ source code. Additionally, when necessary it is also quite easy to customise IOC behaviours on deeper levels, without needing to copy entire IOC source files: eg. the file helpers/st_b2985.py shows how to customise the behaviours of qscan_b2985 in two aspects.
@@
Till now all IOCs discussed in this section interact with hardware other than the controlling computer; besides “sequencers” which will be covered in Section 5, pure-software IOCs can also be written with QueueIOC. Our example for pure-software QueueIOC-based IOCs is qioc_s6, a workalike for the procServControl IOC, based on s6-epics, a workalike of procServ [20]. s6-epics itself is based on s6 (https://skarnet.org/software/s6/), a well-designed suite of programs to manage service processes; that latter most importantly (to us and concerning EPICS) supports separate logging for each service with reliable log rotation to avoid exhaustion of disk space. After organising IOCs in the ~/iocBoot convention, we can use administration commands like those in [2], and a few additional useful commands (Figure 6). As can be seen from the figure, aside from the ability to automatically start specified IOCs upon booting, a currently unique feature of s6-epics is the ability to let both caproto-based IOCs and EPICS IOCs exit gracefully. With qioc_s6, the management of IOCs controlled by s6-epics can be done through the CA protocol; PyDM OPIs (Figure 7) have also been developed for it, available from the supplementary materials.
@@
In EPICS, “sequencers” based on the seq module are used to implement modules like optics, sscan etc; meanwhile, as will be detailed in Section 6, writing seq-based “sequencers” can be error-prone. With QueueIOC the implementation can be much cleaner, by leveraging the expressiveness of Python and following a succinct approach. We would also like to note that as has been hinted in Figure 4(b), a less noticed aspect of “sequencers” is the automated manipulation of PVs in other IOCs through the CA protocol; in our humble opinion this, instead of state transitions, is the real essence of EPICS “sequencers”. Consequently, in QueueIOC although there is a class QSequencerIOC for general state machines (cf. also the qseq_bestec IOC in Section 6), all “sequencers” discussed in this section are actually based on its subclass QMotorSeqIOC. The latter creates state machine with only two states, “down” and “up”, in order to implement a workalike of seq’s “all channels connected & received 1st monitor”. As the name implies, it is mainly used to implement “sequencers” IOCs that (just like those in optics and sscan) manipulate motor IOCs, which we think is a most common application scenario for them. Similar to their EPICS counterparts, all these IOCs expose some motor-like interface, which we have intentionally designed to be unified; PyDM OPIs (Figure 7, available from the supplementary materials) are given for them, along with a Bluesky encapsulation (available as QueueMotor in the open-source edition of Mamba).
@@
The simplest “sequencer” IOCs are used to systematically prevent collision (bumping) between motors; they are based on the class QFuncbumperIOCBase, with qbumper_sim (Figure 8) as a simple example IOC. A more complex QFuncbumperIOCBase-based IOC has been deployed at the transmission X-ray microscope beamline (BE) of HEPS, while more similar IOCs are expected to be deployed at quite a few of the 15 beamlines of HEPS Phase I. As can be seen from qbumper_sim, the complicated logic in the main loop of these “sequencers” IOCs are abstracted in the library, so that developers only need to specify the essential information; in the case of anti-bumping, the information is the mathematical constraints and the list of motor PVs. Under the hood, QFuncbumperIOCBase maintains the state machine in accordance with the underlying motor IOCs; based on this, it denies motion requests when the specified constraints would be violated, when any motor correlated with the requested motor (including itself) is already moving, and when the machine is in the “down” state. Depending on the requirements for other QueueIOC-based “sequencer” IOCs, the information that needs to be specified by the developer can be less mathematical; among them are IOCs based on the class QMotorHubIOCBase, used to control motors connected to multiplexer PLCs (“motor hubs”) for motion controllers. An example IOC based on the class is qmhub_b5, which is actually used in attitude tuning [8] of the Raman spectrometer at the hard X-ray high-resolution spectroscopy beamline (B5) of HEPS, with a simulated variant in the open-source edition of Mamba (cf. also Figure 67). A particularly notable feature of QMotorHubIOCBase is a delay between the latest motion of a motor and the switch from it to any other motor connected to the same multiplexer as this motor, along with another delay after this switch before the newly chosen motor can be moved; both delays are tunable, and are enforced to prevent potential power surges that may damage the devices involved.
@@
Systematic support in QueueIOC for monochromators, a perhaps most widely known application scenario for “sequencer” IOCs, is provided in the class QMonochromatorIOCBase. Based on it, workalikes of the optics IOCs for double-crystal monochromators and high-resolution monochromators are respectively given as qmono_dcm and qmono_hr; a usage example for the latter is given in the supplementary materials. For training purposes, a simple “monochromator” IOC is provided as qmono_sim, which involves two axes and simply uses their sum or difference as the “energy” value. Apart from the basic features, QMonochromatorIOCBase also provides support for automated speed tuning to make the beam change smoothly when the energy value is changed; of course, just like its counterpart in the optics IOC, this class only implements a linear approximation to the ideal behavior. In qmono_dcm and qmono_hr, support is also provided for the user-friendly specification of Miller indices and lattice parameters. This feature, as well as the speed-tuning feature above, are also provided in ways we believe to be most friendly for IOC developers: eg. the latter is enabled by default, and does not need any code to be written by the developer; it can also be opted out by simply overriding the _auto_velo member of QMonochromatorIOCBase. As can be seen from the source code of these IOCs, in our eyes the essence of monochromators IOCs are coordinate transformations between the energy value and the motor positions; calibration is the process of setting up parameters for the transformations before the latter are put into actual use. Therefore QMonochromatorIOCBase can also be used for multiple-to-multiple coordinate transformations: eg. those between the reciprocal and direct spaces in crystallography, where the transformation parameters may be imported from specialised programs like xrayutilities (https://github.com/dkriegner/xrayutilities) and diffcalc (https://github.com/dls-controls/diffcalc). By exposing transformation parameters as PVs, like what is done in the QMonochromatorIOCBase-based IOCs discussed above, the importing process can be automated. Using the PV interface, the parameters can also be exported to other programs: eg. trajectory programs used by motion controllers in variable-speed fly scans involving monochromators or the reciprocal space, which can achieve vastly more accurate motion behaviours than what is possible with the speed-tuning feature of QMonochromatorIOCBase.
6 Case studies: some controllers and “sequencers”
@@
In Section 3, we noted that QueueIOC attempts to replace most EPICS IOCs in a systematic way, while striving to keep the new IOCs as simple as reasonable. We realise that after the brief tour of functionalities in Sections 45, it may still not be obvious what the unique benefits of QueueIOC are; so in this section, we analyse these benefits in detail by having a close look at some example IOCs. Our first example is the qscan_hfda IOC for CNI PSU-H-FDA, PSU-H-LED and PSU-A-D laser controllers. In this IOC, the ancillary class PyHfda and function make_hfda() (Figure 9) are more notable than the IOC class HfdaIOCBase. While the 3 models share a basic serial-based communication protocol, only PSU-A-D supports state readback, and there are also minor differences in the ways to set the laser intensity; the state readback (cf. the function read_state() in PyHfda) provides more than 10 information items, all of which are mapped into PVs by the IOC. The state-changing commands require checksum bytes and produce echo replies (the models also differ slightly in their echo behaviours, as can be seen from send_cmd() in PyHfda), so in case a setpoint item is not available from the state readback, the IOC needs to decide whether to update the value of the corresponding PV upon user writing based on the correctness of the replies to the commands. For this kind of readback-less setpoints (enable_val, as well as current_val and power_val for models other than PSU-A-D), the IOC also needs to properly initialise the corresponding hardware settings, so that the initial values of their PVs are correctly reflected. Moreover, the controllers may encounter read timeouts if a state-reading command is sent too closely after a state-changing command, so an additional delay should be enforced after the latter. All the above are handled succinctly and cleanly in qscan_hfda: state readback in read_state() based on a call to struct.unpack(); complete correctness check in send_cmd() and read_state() based on checksums and echo replies; the automatic differentiation between the models, as well as the preparation of initial states, in make_hfda(). Done in HfdaIOCBase are the automatic construction of PV lists for different models, the different handling of state changing depending on the model, and the delay after state changing. While these are not impossible in EPICS IOCs, they would be very awkward to implement: what the struct.unpack() does above would be highly bloated with StreamDevice; additionally StreamDevice cannot easily support the echo check and the delay after state changing, and we also do not find seq-based “sequencers” very helpful on this issue. Writing the IOC in C/C++ may be also considered, but the codebase would still be significantly larger, especially considering the support for different versions; these make the development and maintenance cost of such an IOC much higher than our IOC.
@@
Admittedly, with caproto, PCASPy etc, it would not be hard to write workalikes of qscan_hfda that are not much more complex than it. This is because its logic revolves around simple state-changing and state-reading commands, and the abstraction for this pattern, QScanIOC, is easy to reimplement; as has been hinted above, PyHfda and make_hfda() are what really make the IOC simple. There are also examples where the tools provided by QueueIOC significantly contribute to the simplicity of the IOCs: in fact, all IOCs in Section 5 can be seen as examples for this; it can be easily seen when the reader considers how the same requirements could be implemented with EPICS, caproto, PCASPy etc. Another example is the “sequencer” IOC qseq_bestec available from the supplementary materials, used with a Bestec optical complex (PGM, slit & mirror unit) at the high-resolution nanoscale electronic structure spectroscopy beamline (BC) of HEPS. The vendor supplies a proprietary IOC-like program that encapsulate the device interfaces as waveform PVs, which are essentially argument lists for control commands. As what we really want are scalar PVs, this IOC was written to translate the argument lists in “master” PVs into single elements in “slave” PVs, and translate caput operations on slave PVs back into caput on the master PVs. In this IOC, a table of master-PVs and their slave-PVs is first created; then accordingly, the slave PVs are created and the master PVs are connected to; in the main event loop, caput operation on the slave PVs and monitored updates of the master PVs are handled. This IOC is simple and easily adaptable to another set of master and slave PVs; this is because all the caput and monitor events are represented in a well-designed way there, allowing for a single function to handle all the PVs without repeated code. The similar requirements would be much less convenient to implement, whether with caproto, PCASPy etc or (especially) with EPICS and seq.
@@
In our summary, what distinguishes QueueIOC from the others is an intentional and systematic pursuit of the utmost simplicity [21], or in other words approaching the complexity/cost lower-bounds (cf. Section 1); what QueueIOC attempts to do, in practice, is to provide succinct yet powerful tools that developers can use to build IOCs that are close to their complexity/cost lower-bounds. Here we also discuss monochromators in detail, as another practical demonstration of how simplicity is achieved in QueueIOC-based IOCs – with not only particular tools, but also reusable design patterns. Writing seq-based “sequencers” is often error-prone, because the state transitions in them can easily become comparable to the “goto hell” C/C++ (Figure 10c), result in hard-to-trace bugs (a few of them are given in the supplementary materials). This is why most “sequencers” in QueueIOC only have the “down” and “up” states (cf. Section 5); the domain-specific business logic is instead encapsulated in the state-transition functions. In monochromator IOCs based on QMonochromatorIOCBase, the functions serve_down(), serve_up(), do_common() etc are the basis of our abstraction of their business logic as coordinate transformations. However, in addition to these shared logic, real-world monochromators need invertible three-way conversions between the Bragg angle $\theta$, the wavelength $\lambda$ and the energy $E$; the conversions can be even more complex for certain monochromators, eg. the high-resolution monochromator (Figure 10a). To do this cleanly, we use functions to encode these conversions: eg. mono_phis() in the qmono_hr IOC (Figure 10b), as well as mono_theta() which is used by qmono_dcm and usable by other simple variants of the double-crystal monochromator.
@@
In comparison with the monochromator IOCs in optics, IOCs like qmono_dcm are not only much shorter, but also much better modularised. The shared logic is contained in the code of QMonochromatorIOCBase, organised cleanly unlike the “goto hell”. The $\theta$$\lambda$/$E$ conversions are fully encapsulated in functions like mono_phis(), which are implemented as straightforward encodings of graphs like Figure 10(a). The rest cannot be easily abstracted as libraries, but they are still organised cleanly, with a structure intentionally kept consistent across different monochromator IOCs; this can be seen from a comparison between the IOC classes, eg. MonoHrIOC in qmono_hr and MonoDcmIOC in qmono_dcm. The result of the structural differences above is dramatic reduction in the cost of development and maintenance. When writing a new monochromator IOC, the developer often only needs to customise the conversion function and the IOC class, without worrying about the possibility of a small change conflicting with some assumptions implicit in the global state. When debugging, issues can usually be easily traced and resolved thanks to the careful decoupling: eg. the conversion function can isolated and used as a calculator for the variables involved, greatly facilitating tests. The same approach is also followed elsewhere in QueueIOC for non-trivial IOCs: aside from PyHfda in qscan_hfda, another example is the qscan_redpower IOC for SPI redPOWER laser controllers based on helpers/qdev_redpower_lib.py, a self-made mini-“SDK”. In the complex feature set provided by the controllers, this SDK covers a non-trivial subset currently demanded by the high-pressure beamline (B6) of HEPS, where those controllers are used. Similarly, in QDetectorIOC [19], vendors’ SDKs are also encapsulated thinly into mini-SDKs; many of these SDKs can dynamically adapt to variable device interfaces. These self-made SDKs expose succinct, reusable and adaptable interfaces, which can be tested in isolation and even reused in standalone applications. In summary, modularity facilitates reuse, which enhances regularity and therefore helps to reduce the complexity of the entire system (cf. also the discussion on PVA, TANGO and Karabo in Section 3). With measures taken like those above, the IOCs provided by QueueIOC have been made satisfactorily close to their complexity/cost lower-bounds. As has been discussed in Section 1, with the consistent pursuit of simplicity throughout QueueIOC, we believe it is able to bring about significant efficiency boosts in the EPICS ecosystem, and help large scientific facilities (along with their staffs) become more intelligent.